The distance between the centres of carbon and oxygen atoms in carbon monoixide molecule is 0.12nm  .What will be the distance of tbe centre if mass of the molecule from the carbon atom .?

Grade:

1 Answers

Arun
5 years ago
The distance between carbon and oxygen atoms in CO is 1.2 10-10 m (= 1.2A0).
 
Given, distance between the atoms (x2) = 1.2 A0
Mass of carbon atom (m1) = 12
Mass of oxygen atom (m2) = 16
Consider, the situation when C-atom is located at origin, it is assumed this way because, center of mass has to be calculated with respect to C-atom.
 
Xcm = 12*0 + 16 * 1.2 / 28 = 19.2/28 = 0.68 A